A1,A2,A3.......A10 ARE IN AP H1,H2,H3....H10 ARE IN H.P IF A1=H1=2,A10=H10=3 THEN A4*H7 IS??

Grade:11

2 Answers

shubham sharma
10 years ago

Its 6

first find common differences from given values and then a4 &h7 !!!

Khandavalli Satya Srikanth
10 years ago

nth term in AP=a+(n-1)d

3=2+(10-1)d

d=1/9

A4=2+(4-1)*1/9=7/3

similarly (1/3)=(1/2)+(10-1)d (since if a,b,c are in HP,1/a,1/b,1/c are in AP)

d=(-1/54)

H7=[1/{(1/2)+(7-1)*(-1/54)}]

H7=2.5714

A4*H7=(7/3)*(2.5714)=6